OpenStudy (anonymous):

Find the 10th term of the sequence 64, -32, 16, -8, …

OpenStudy (anonymous):

a=64 r=-1/2 10term= ar^n-1 =64(-1/2)^9 -2^6(2^-9)=-1/8

OpenStudy (anonymous):

For real nice

OpenStudy (anonymous):

any time! I hope u understand my method of writing.